Journalist Report – February 22th

Sol 10 – Wuthering Heights and Rainy Nights

“The rockets came like locusts, swarming and settling in blooms of rosy smoke.”

– – Chapter 10 of The Martian Chronicles by Ray Bradbury

Yesterday, I wrote that Mars had spared us from rain… As though summoned, it fell during the night between Sols 9 and 10. Quentin, our crew engineer, the only crew member sleeping in a mezzanine above the Upper Deck, was woken up at 4 in the morning: “The Martian winds were so strong, I thought the Hab roof would be blown away. It was extremely noisy, and instinctively I climbed down the ladder to sleep on the Upper Deck.”

In the morning, a single question was on everyone’s mind: would we be able to go out on EVA to retrieve the LOAC, which we feared had been damaged by the rain, and execute the second phase of the photogrammetry experiment? During breakfast, after our daily workout, we let out a sigh of relief: Mission Support cleared us to exit the Hab and not only retrieve the atmospheric equipment, but to go to North Ridge as planned and conduct the photogrammetry EVA. This would be my first outing as EVA leader.

After depressurizing the airlock, Alice, Corentin and I stepped outside of the Hab, exposing ourselves to the Martian winds. Upon arriving at the MegaAres site, we contacted HabCom to bitterly report that the weather station and LOAC had been toppled by the wind. The PurpleAir atmospheric sensor was also detached from the weather station mast, and mud had infiltrated the electronics compartment. It was decided on the spot to retrieve this instrument as well as the LOAC, which is still operational! The MegaAres mast is also still in its upright position. According to HabCom, a cry of victory resonated in the Upper Deck when I passed on the news.

We then proceeded to North Ridge to begin the second exploration EVA for the photogrammetry experiment. Upon arrival, we noticed the winds were much stronger, accelerating in narrow gullies. This made it very difficult to communicate, with the sound of the wind resonating around our suits and effectively drowning out our radio communications. Radio contact with HabCom was also very choppy and faint. After about 45 minutes on site and having reached 3 of the 6 checkpoints, the wind had become so strong that we were starting to lose balance. We found temporary cover, and after trying to contact the Hab, I decided to terminate the EVA and return for the safety of the crew. I found that removing my spacesuit and warming up besides the Lower Deck heater after a trying outing was the best feeling in the world… After debriefing and drinking tea flavored with fresh mint leaves from the GreenHab, we stayed gathered around the Hab table to simply talk and debate amongst ourselves, taking a necessary break from “MDRS talk”.

By early afternoon, the winds had not waned, effectively cancelling the second EVA of the day, the purpose of which being to reinstall the repaired MegaAres antenna. The whirling and thunder-y sound of the wind was almost inebriating as I stayed in the Upper Deck most of the afternoon.

Some news from the GreenHab: the aquaponic system is nominal, all fishes are still alive and well, and the plants are steadily growing. Adrien is performing daily tests on the water, checking for excesses of any substances that could deregulate the system and affect the health of the plants and fishes.

Quentin is also proud to announce that all environmental sensors are deployed within the station and operational! They will provide additional data for our human factors experiments, and be used to test an AI developed by CNES, SPooN, and students from ISAE-Supaero. AI4U, whose protocols Alexandre has started testing in the last few days, is an AI designed to assist astronauts in their tasks, which we will specifically use to assist us during mock emergency protocols.

All crew members having gathered around the dining table for the Comms window, we laughed as we speculated on what Mars could throw at us next… clearly, he still has surprises in store for us!

EVA Report – February 22th

Crew 275 EVA Report 22Feb2023

EVA # 8

Author: Marie Delaroche, EVA Leader and Crew Journalist

Purpose of EVA: The purpose of the EVA was to perform the photogrammetry experiment with the 2D map.

First, as we were on our way to North Ridge, we checked on the LOAC, the Purple Air instrument and the MegaAres.

We then went to North Ridge where the two subjects of the experiment tried to reach specific checkpoints while the operator of the experiment onsite took time measurements and drew their trajectory on a map. In addition, the two subjects were equipped with a step counter.

Note: Because it rained during the night and there was further risk of precipitation, we focused on retrieving the LOAC, as this device can be damaged by rain and humidity.

Start time: 9:10

End time: 10:30


09:10: Airlock closed.

09:15: EVA crew exits airlock.

09:17: All equipment stowed in Perseverance trunk, both rovers unplugged.

09:19: EVA crew is cleared for departure by HabCom.

09:21: Arrival at the MegaAres site near Marble Ritual. Due to the wind, we found that the weather station has collapsed. Sensitive equipment on the mast is secured, no attempts to put it upright are attempted. The MegaAres mast is still upright and LOAC is still functional.

09:32: Return to rovers, carrying the LOAC in its case with its battery, both loaded into Perseverance’s trunk.

09:35: Departure from the location of the instruments, between the Hab and Marble Ritual.

09:38: Arrival at North Ridge, radio contact with HabCom through Channel 1 is often cut off.

09:41: EVA crew begins exploring the area to reach pre-defined checkpoints.

10:00: Contact between EVA leader and EVA crew is partially lost, EVA leader switches to Channel 2 to hear her teammates better. The wind makes it difficult to hear each other.

10:08: Three out of the six checkpoints are reached, but the wind is becoming stronger and EVA crewmembers are cold. Decision to shorten EVA and return to the Hab for the crew’s safety.

10:18: EVA crew is back at the rovers.

10:20: Departure from North Ridge

10:24: EVA crew is back at the Hab, rovers are parked and plugged in.

10:30: EVA crew is safely inside the Hab.

Destination: North Ridge (first, we stopped where the MegaAres and the weather station are located)

Coordinates: N4252000 E519000

Participants: EVA Leader: Marie Delaroche (Crew Journalist), Alice Chapiron (Crew Scientist), and Corentin Senaux (HSO)

Road(s) and routes per MDRS Map: Cow Dung Road until North Ridge, near the intersection with “Gateway to Candor”. we then walked North-West, up to an estimated 300 meters from the rover.

Mode of travel: Driving using Perseverance and Opportunity

Crew 275 EVA Report 22-02-2023

EVA # 9

Author: Alexandre VINAS (crew astronomer and EVA leader)

Purpose of EVA: Installation of the LOAC instrument & Maintenance of the other atmospheric instruments

Start time: NA (planned 14:00)

End time: NA (planned 17:00)


EVA canceled because of too high winds

Destination: /
Coordinates (use UTM WGS 84): NA (planned N4251000 E518500)

Participants: EVA leader : Alexandre Vinas (Crew Astronomer), Marie Delaroche (Crew Journalist), Corentin Senaux (HSO)

Road(s) and routes per MDRS Map: NA (planned drive Cow Dung Road 0110 to a place just before Marble Ritual (see map on the attached picture))

Mode of travel: NA (planned Spirit and Curiosity)

GreenHab Report – February 22th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 65.4 °F

Average humidity: 31 %

Hours of supplemental light: 0 hour

Daily water usage for crops: 10.57 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 76.0 gallons

Time(s) of watering for crops: 9:24 & 14:47

Changes to crops: N/A

Narrative: This morning I watered and maintained crops and plants as my daily activities.

There were some dry leaves on some tomato plants, but overall, everything looks great. Plants have grown well and look good.

The dry leaves are located just where the heater hit, so we can assume that it is the heater that dries the leaves.

What I harvested allowed us to cook a nice meal, beginning with a couscous and finishing with a nice mint tea!

Harvest: 65g of beans, 3g of mint, and 2g of thyme

Support/supplies needed: N/A

Operations Report – February 22th

Crew 275 Operations Report 22Feb2023
SOL: 10

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9

Notes on non-nominal systems:
• See section “Summary of suits and radios” below

Spirit rover used: No
• Hours: 209.1
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: Yes
• Hours: 113.2
• Beginning charge: 100%
• Ending charge: 70%
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: Yes
• Hours: 255.6
• Beginning charge: 100%
• Ending charge: 61%
• Currently charging: Yes

General notes on rovers: Rovers were used during today’s EVA. Nothing to report.

Summary of Hab operations: The Crew conducted routine operations. Due to high winds, the main Airlock door and the Engineering Airlock doors were closed and secured.
• WATER USE: 13.90 gallons
• Water: 329.08 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS.

Summary of suits and radios: During EVA preparation, as a Crew Member was checking the battery of the Spacesuit n°9, he accidentally opened too high the back wooden panel. The panel was subsequently broken, and the panel separated from the spacesuit, breaking at the hinges. I can provide additional photos if you need.

Summary of GreenHab operations: The Crew botanist conducted routine operations.
• WATER USE: 10.57 gallons
• Heater: On
• Supplemental light: No
• Harvest: 65 g of beans – 3 g of mint – 2 g of thyme

Summary of ScienceDome operations: The Crew Scientist conducted routine operations. The preparation of the EVA concerning the photogrammetry experiment was performed.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Today’s EVA was aborted after 80 minutes on field, because of the very high winds and the cold temperature felt at North Ridge.

Questions, concerns, and requests to Mission Support:
• Would it be possible to have a supplemental table and a supplemental chair in the GreenHab? Indeed, we need to perform psychometric tests there.

Sol Summary Report – February 22th

Crew 275 Sol Summary Report 22Feb2023

Sol: 10

Summary Title: Wuthering Heights and Rainy Nights

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Since yesterday evening, our activities have been highly impacted by the weather conditions. The noise of the wind was very impressive and impaired the sleep of some crewmembers. We were all very afraid for our atmospheric experiments, especially for the LOAC instrument that is very sensitive to rain. In the morning, we were authorized to leave the Hab for EVA #8 and stopped on our way to check the status of our instruments, as per EVA request. We noticed that both the meteorological mast and the LOAC had toppled over. We only retrieved the LOAC and left the meteorological mast, because we were aware that the wind speed would remain quite high. The crew of EVA #8 then reached North Ridge to perform their main mission, but after about 40 minutes they decided to come back because they were too cold, and the wind speed was too high to safely conduct their operations.

Following this event, we decided to also cancel EVA #9 that was planned for the afternoon.

On the campus, the environmental sensors have been fully deployed on all the buildings and can now give us information about temperature, humidity, pressure, and luminosity.

Look Ahead Plan:

We had to adapt our plans because of the weather. We will wait for better wind conditions to put back the antenna of Mega-Ares back at its initial place.

Today, two crewmembers have trained with the 3D map of North Ridge obtained by photogrammetry. Tomorrow, we would like to test their ability to reach some specific checkpoints there, as compared to the EVA crew of today (using only a 2D map). We are thus planning an EVA tomorrow morning (Sol 11), should the weather conditions allow it.

Anomalies in work: None.

Weather: Very cloudy and very windy,low 0°C/32°F, high 9°C/48°F.

Crew Physical Status: Optimal.

EVA: Because of the bad weather conditions, the team of EVA#8 had to retrieve the LOAC instrument, which was done successfully. They had to shorten their EVA because of high wind, which means that the scientific objectives were only partially achieved. EVA #9 was cancelled because the LOAC instrument was already retrieved, and the weather conditions were not good enough to safely leave the Hab.

Reports to be filed: Sol Summary, EVA Request, EVA Report, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site